User’s Guide to the autofilo package^{∗}

Frank G. Bennett, Jr.

December 9, 1998

Abstract This package provides an environment within which pages are framed with cut lines and printed with punch-marks, so that printed text can easily be inserted into a filofax or binder.

1 About the Style

This style provides a rough solution to the design problem it addresses, but it does work. Pages are always printed in landscape, rotated using the lscape package from the graphics bundle, and centered on the physical paper.^{1} It accepts two options, twocolumn (the default) and fourcolumn.

The style provides an environment called autofilo. Most often, users of this package will be style authors who want to incorporate it into other environments that do more specific things. Because autofilo accepts several keyval-style parameters, style authors may have a use for documentation for those optional parameters. So here is some sample documentation:

punchcluster Most filofaxes have two or or more clusters or groups of punches to hold the pages in place. This option sets the number of punches in each cluster. The default is three.

intraspace This governs the space between punchouts within a group. The default is 19.25mm.

punchgroups This option sets the number of groups of punches. The default is two groups.

interspace This option sets the distance between the groups of punches. The default is 51.25mm.

pageheight This fixes the height of a filofax page (not the physical paper on which it is printed). The default is 172mm.

pagewidth This fixes the width of an individual page. The physical printed area will be twice this figure. The default is 95mm, for a 190mm printed area. grip This adjusts the distance from the edge of the page to the outer edge of the

∗This file is version number 3.1.1.1. It was last revised on 1998/01/17 18:13:42.

1As a technical note, the frame itself is not rotated, because it is inserted independently of the output boxes controlled by the L^{A}TEX output routine. The output modes (two-column, and a hacked-together four-column mode) have been modified to understand blocks of text as \hboxes to be laid atop one another, rather than as \vboxes to be set alongside one another.

punchouts. Defaults to 5mm.

punchmargin This adjusts the distance from the edge of the text to the inner edge of the punchouts. Defaults to 2mm.

punchsize Size, in points, of punchouts. Defaults to 15.

topspace Gap between top of filofax page and top of text page.

bottomspace Gap between bottom of filofax page and bottom of text page.

jawspread If set to a positive length, this places a set of rules the width of the punchmarks on either side of each punchhole, centered on its center and spread the distance specified. This can be useful as a guide with some onehole punches that are designed to be used “blind”.

jawline Sets the width of the lines used to make jawmarks. The default value is

     0.4pt.